Homework: HW5 Problem 4-77 (algorithmic) a. If F $9,500, G-$550, and N-6, then i? d.lfFe$8,000, N-6, and i-8% per period, then G#7 to one
Answer #1

Answer

a) pv of anuity F =G*(1-(1+i)^-6)/i

9500=550*(1-(1+i)^-6)/i

So solving for i we get i=23.812%

B) pv of annuity F=G*(1-(1+i)^-n) /i

9500=550*(1-(1+4%)^-n)/4%

n=30 years

C)

Pv of annuity F=1200*(1-(1+8%)^-12)/8%

F=$9043.29

D) Annuity G=F*i/(1-(1+i)^-n)

G=8000*8%/(1-(1+8%)^-6)

G=$1730.52